# Improved Particle Approximation Error for Mean Field Neural Networks

### v2

@@ Introduction
where $F:\mathcal{P}_{2}(\mathbb{R}^{d})\rightarrow\mathbb{R}$ is a convex functional on the space of probability distributions $\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\mathrm{Ent}(\mu)=\int\mu(\mathrm{d}x)\log\frac{\mathrm{d}\mu}{\mathrm{d}x}(x)$ is a negative entropy. Especially we focus on the learning problem of mean-field neural networks, that is, $F(\mu)$ is a risk of (infinitely wide) two-layer neural networks $\mathbb{E}_{X\sim\mu}[h(X,\cdot)]$ where $h(X,\cdot)$ represents a single neuron with parameter $X$. One advantage of this problem is that the convexity of $F$ with respect to $\mu$ can be leveraged to analyze gradient-based methods for a finite-size two-layer neural network: $\frac{1}{N}\sum_{i=1}^{N}h(x^{i},\cdot)$ ($x^{i}\in\mathbb{R}^{d}$). This is achieved by translating the optimization dynamics of the finite-dimensional parameters $(x^{1},\ldots,x^{N})\in\mathbb{R}^{dN}$ into the dynamics of $\mu$ via the mean-field limit: $\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{i}}\rightarrow\mu~{}(N\rightarrow\infty)$. This connection was pointed out by Nitanda and Suzuki, (2017); Mei et al., (2018); Chizat and Bach, (2018); Rotskoff and Vanden-Eijnden, (2022) in the case of $\lambda=0$, and used for showing the global convergence of the gradient flow for (1) by Mei et al., (2018); Chizat and Bach, (2018).
@@
where $\{W_{t}\}_{t\geq 0}$ is the $d$-dimensional standard Brownian motion and $\frac{\delta F}{\delta\mu}(\mu)$ is a first-variation of $F$. Even though several optimization methods (Nitanda et al.,, 2021; Oko et al.,, 2022; Chen et al.,, 2023) that can efficiently solve the above problem with polynomial computational complexity have been proposed, MFLD remains an interesting research subject because of the above connection to the noisy gradient descent. In fact, recent studies showed that MFLD globally converges to the optimal solution (Hu et al.,, 2019; Mei et al.,, 2018) thanks to noise perturbation and that its convergence rate is exponential in continuous-time under uniform log-Sobolev inequality (Nitanda et al.,, 2022; Chizat,, 2022).
@@ 1.1 Contributions
Here, we mention the proof strategy of Theorem 1. Langevin dynamics (LD) is a special case of MFLD corresponding to the case where $F$ is a linear functional. It is well known that even with a single particle, we can simulate LD and the particle converges to the target Gibbs distribution. This means that the particle approximation error is due to the non-linearity of $F$. Therefore, in our analysis, we carefully treat the non-linearity of $F$ and obtain an expression for the particle approximation error using the Bregman divergence induced by $F$. Finally, we reduce the evaluation of this divergence to the application of concentration inequalities by utilizing the problem structure. This proof strategy is quite different from existing ones and is simple. Moreover, it leads to an improved approximation error as mentioned above. We refer the readers to Section 4 for details about the proof.
@@ 2.1 Problem setting
and say $G$ is convex when it follows that for $\forall\mu,\mu^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
@@
There exists $M_{1},M_{2}>0$ such that for any $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $x\in\mathbb{R}^{d}$, $\left|\frac{\delta F_{0}(\mu)}{\delta\mu}(x)\right|<\infty$, $\left|\nabla\frac{\delta F_{0}(\mu)}{\delta\mu}(x)\right|\leq M_{1}$ and for any $\mu,\mu^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $x,x^{\prime}\in\mathbb{R}^{d}$,
@@ 2.2 Mean-field Langevin dynamics and finite-particle approximation
where $\xi_{k}^{i}\sim\mathcal{N}(0,I_{d})~{}(i\in\{1,\ldots,d\})$ are i.i.d. standard normal random variables. By extending the proof techniques developed by Chen et al., (2022), Suzuki et al., 2023b proved the uniform-in-time propagation of chaos for MFLD (12); there exist constants $C_{1},C_{2}>0$ such that
@@
where $\mu^{(N)}_{k}=\mathrm{Law}(\mathbf{X}_{k})$. The last two terms are due to time-discretization and finite-particle approximation, respectively. The finite-particle approximation error $O(\frac{\lambda}{\alpha N})$ appearing in both continuous- and discrete-time settings (11), (13) means the deterioration as $\alpha\rightarrow 0$. Considering typical estimation $\alpha\gtrsim\exp(-\Theta(1/\lambda))$ (e.g., Theorem 1 in Suzuki et al., 2023b ) of LSI-constant using Holley and Stroock argument (Holley and Stroock,, 1987) or Miclo’s trick (Bardet et al.,, 2018), these bounds imply that the required number of particles increases exponentially as $\lambda\rightarrow 0$.
@@ 3.1 LSI-constant free particle approximation error for mean-field neural networks
For $\forall y\in\mathcal{Y}$, $\ell(\cdot,y)$ is convex and Lipschitz smooth, and $h(\cdot,\cdot)$ is uniformly bounded;
@@
  • There exists $R>0$ such that for any $x\in\mathbb{R}^{d},z\in\mathcal{Z}$, $|h(x,z)|\leq R$.
We easily verify this assumption for mean-field neural networks using a bounded activation function (Nitanda et al.,, 2022; Chizat,, 2022; Chen et al.,, 2022; Suzuki et al., 2023b, ) and standard loss functions such as logistic loss and squared loss. The following is the main theorem that bounds $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})-\mathcal{L}(\mu_{*})$. The proof is deferred to Section 4 and Appendix 5.1.
@@
$\frac{\lambda}{N}\mathrm{KL}(\mu^{(N)}_{*}\|\mu^{\otimes N}_{*})\leq\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})-\mathcal{L}(\mu_{*})\leq\frac{2LR^{2}}{N}.$ (15)
@@ 3.2.1 Convergence of the mean-field Langevin dynamics
By setting $g=\sqrt{\frac{\mathrm{d}\mu^{(N)}}{\mathrm{d}\mu^{(N)}_{*}}}$, Assumption 4 leads to $\mathrm{KL}(\mu^{(N)}\|\mu^{(N)}_{*})\leq\frac{1}{2\bar{\alpha}}\mathbb{E}_{\mu^{(N)}}\left[\left\|\nabla\log\frac{\mathrm{d}\mu^{(N)}}{\mathrm{d}\mu^{(N)}_{*}}\right\|_{2}^{2}\right]$. Using the tensorization argument (Proposition 5.2.7 in Bakry et al., (2014)) with Lipschitz perturbations (Cattiaux and Guillin,, 2022), Kook et al., (2024) established LSI on $\mu^{(N)}_{*}$ with a constant independent of $N$ as follows. Under the boundedness condition (Assumption 1): $\left|\nabla\frac{\delta F_{0}(\mu)}{\delta\mu}(x)\right|\leq M_{1}$, $\mu^{(N)}_{*}$ satisfies LSI with a constant $\bar{\alpha}$ given below (Lemma 22 in Kook et al., (2024)):
$\bar{\alpha}\geq\frac{\lambda^{\prime}}{8\lambda}\wedge\frac{1}{2}\left(\ \frac{M_{1}}{\lambda^{\prime}}+\sqrt{\frac{\lambda}{\lambda^{\prime}}}\right)^{-2}\left(2+d+\frac{4M_{1}^{2}}{\lambda^{\prime}\lambda}\right)^{-1}\exp\left(-\frac{M_{1}^{2}}{2\lambda^{\prime}\lambda}\right).$
@@
    $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{t})-\mathcal{L}(\mu_{*})\leq\frac{2LR^{2}}{N}+\exp(-2\bar{\alpha}\lambda\eta t)\left(\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{0})-\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})\right),$
@@
    $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{k})-\mathcal{L}(\mu_{*})\leq\frac{2LR^{2}}{N}+\frac{\delta^{(N)}_{\eta}}{2\bar{\alpha}\lambda}+\exp(-\bar{\alpha}\lambda\eta k)\left(\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{0})-\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})\right),$
@@
The term of $\frac{2LR^{2}}{N}$ is the particle approximation error inherited from Theorem 1. Again our result shows the LSI-constant independence particle approximation error for MFLD unlike existing results (Chen et al.,, 2022; Suzuki et al., 2023b, ) where their error bounds $O(\frac{\lambda}{\alpha N})$ scale inversely with LSI-constant $\alpha$ as seen in (11) and (13). Hence, the required number of particles to achieve $\epsilon$-accurate optimization: $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)})-\mathcal{L}(\mu_{*})\leq\epsilon$ suggested by our result and Chen et al., (2022); Suzuki et al., 2023b are $N=O(\frac{1}{\epsilon})$ and $N=O(\frac{\lambda}{\alpha\epsilon})$, respectively. Whereas the iterations complexity of MFLD (12) is $O(\frac{1}{\bar{\alpha}^{2}\lambda\epsilon}\log\frac{1}{\epsilon})$ which is same as that in Suzuki et al., 2023b up to a difference in LSI constants $\alpha$ or $\bar{\alpha}$.
@@ 3.2.2 Sampling guarantee for $\mu_{*}$
    $\lambda\mathrm{KL}(\mu^{(N)}_{t,1}\|\mu_{*})\leq\mathcal{L}(\mu^{(N)}_{t,1})-\mathcal{L}(\mu_{*})\leq\frac{2LR^{2}}{N}+\exp(-2\bar{\alpha}\lambda\eta t)\Delta^{(N)}_{0},$
@@
    $\lambda\mathrm{KL}(\mu^{(N)}_{k,1}\|\mu_{*})\leq\mathcal{L}(\mu^{(N)}_{k,1})-\mathcal{L}(\mu_{*})\leq\frac{2LR^{2}}{N}+\frac{\delta^{(N)}_{\eta}}{2\bar{\alpha}\lambda}+\exp(-\bar{\alpha}\lambda\eta k)\Delta^{(N)}_{0}.$
@@
For any distribution $\mu^{(N)}\in\mathcal{P}_{2}(\mathbb{R}^{dN})$ whose marginal $\mu^{(N)}_{i}$ on $i$-th coordinate $x^{i}$ ($i\in\{1,\ldots,N\}$) are identical each other, it follows that by the convexity of the objective function and the entropy sandwich (Nitanda et al.,, 2022; Chizat,, 2022): $\lambda\mathrm{KL}(\mu\|\mu_{*})\leq\mathcal{L}(\mu)-\mathcal{L}(\mu_{*})$ $(\forall\mu\in\mathcal{P}_{2}(\mathbb{R}^{d}))$,
@@ 3.2.3 Uniform-in-time Wasserstein propagation of chaos
    $\frac{1}{N}W_{2}^{2}(\mu^{(N)}_{t},\mu^{\otimes N}_{t})\leq\frac{4}{\alpha\lambda}\left(\frac{2LR^{2}}{N}+\exp(-2\bar{\alpha}\lambda\eta t)\Delta^{(N)}_{0}+\exp(-2\alpha\lambda t)\Delta_{0}\right).$
@@
    $\frac{1}{N}W_{2}^{2}(\mu^{(N)}_{k},\mu^{\otimes N}_{k})\leq\frac{4}{\alpha\lambda}\left(\frac{2LR^{2}}{N}+\frac{\delta^{(N)}_{\eta}}{2\bar{\alpha}\lambda}+\frac{\delta_{\eta}}{2\alpha\lambda}+\exp(-\bar{\alpha}\lambda\eta k)\Delta^{(N)}_{0}+\exp(-2\alpha\lambda k)\Delta_{0}\right),$
@@
Applying the convergence rates of finite- and infinite-particle MFLDs (Nitanda et al., (2022) and Theorem 2), we conclude the proof. For completeness, we include the auxiliary results used in the proof in Appendix 6. ∎
Corollary 2 uniformly controls the gap between $N$-particle system $\mu^{(N)}_{t}/\mu^{(N)}_{k}$ and its mean-field limit $\mu^{\otimes N}_{t}/\mu^{\otimes N}_{k}$. Again this result shows an improved particle approximation error $O(\frac{1}{\alpha\lambda N})$ over $O(\frac{1}{\alpha^{2}N})$ (Chen et al.,, 2022; Suzuki et al., 2023b, ).
@@
In this section, we provide the proof sketch of Theorem 1. Our analysis carefully treats the non-linearity of $F_{0}$ because the differences between $\mathcal{L}$ and $\mathcal{L}^{(N)}$, and between their solutions $\mu_{*}$ and $\mu^{(N)}_{*}$ come from this non-linearity. In fact if $F_{0}$ is a linear functional: $F_{0}(\mu)=\mathbb{E}_{\mu}[f]$ ($\exists f:\mathbb{R}^{d}\rightarrow\mathbb{R}$), then $\mathcal{L}^{(N)}(\mu^{(N)})=\sum_{i=1}^{N}\mathbb{E}_{X^{i}\sim\mu^{(N)}_{i}}[f(X^{i})+\lambda^{\prime}\|X^{i}\|_{2}^{2}]+\lambda\mathrm{Ent}(\mu^{(N)})\geq\sum_{i=1}^{N}\mathcal{L}(\mu_{i}^{(N)})$, where $\mu^{(N)}_{i}$ are marginal distributions on $X^{i}$. This results in $\mu^{(N)}_{*}=\mu^{\otimes N}_{*}$ and $\mathcal{L}^{(N)}(\mu^{(N)}_{*})=N\mathcal{L}(\mu_{*})$, and thus there is no approximation error by using finite-particles as also deduced from Theorem 1 with $L=0$. Therefore, we should take into account the non-linearity of $F_{0}$ to tightly evaluate the gap between $\mathcal{L}$ and $\mathcal{L}^{(N)}$.
@@
$=\exp\left(-\frac{N}{\lambda}\left(F(\mu)+\left\langle\frac{\delta F}{\delta\mu}(\mu),\mu_{\mathbf{x}}-\mu\right\rangle+B_{F}(\mu_{\mathbf{x}},\mu)\right)\right)$
$\propto\exp\left(-\frac{N}{\lambda}B_{F}(\mu_{\mathbf{x}},\mu)\right)\prod_{i=1}^{N}\exp\left(-\frac{1}{\lambda}\frac{\delta F}{\delta\mu}(\mu)(x^{i})\right)$
@@
Intuitively, $\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[B_{F}(\mu_{\mathbf{X}},\mu_{*})\right]$ is small because the empirical distribution $\mu_{\mathbf{X}}$ ($\mathbf{X}\sim\mu^{\otimes N}_{*}$) converges to $\mu_{*}$ by law of large numbers. In fact, we can reduce the evaluation of this term to the application of a concentration inequality, yielding a bound: $\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[B_{F}(\mu_{\mathbf{X}},\mu_{*})\right]\leq\frac{2LR^{2}}{N}$ (see the proof of Theorem 1 in Appendix 5). Then, we arrive at Theorem 1.

### v3

@@ Introduction
where $F:\mathcal{P}_{2}(\mathbb{R}^{d})\rightarrow\mathbb{R}$ is a convex functional on the space of probability distributions $\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\mathrm{Ent}(\mu)=\int\mu(\mathrm{d}x)\log\frac{\mathrm{d}\mu}{\mathrm{d}x}(x)$ is a negative entropy. Especially we focus on the learning problem of mean-field neural networks, that is, $F(\mu)$ is a risk of (infinitely wide) two-layer neural networks $\mathbb{E}_{X\sim\mu}[h(X,\cdot)]$ where $h(X,\cdot)$ represents a single neuron with parameter $X$. One advantage of this problem is that the convexity of $F$ with respect to $\mu$ can be leveraged to analyze gradient-based methods for a finite-size two-layer neural network: $\frac{1}{N}\sum_{i=1}^{N}h(x^{i},\cdot)$ ($x^{i}\in\mathbb{R}^{d}$). This is achieved by translating the optimization dynamics of the finite-dimensional parameters $(x^{1},\ldots,x^{N})\in\mathbb{R}^{dN}$ into the dynamics of $\mu$ via the mean-field limit: $\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{i}}\rightarrow\mu~{}(N\rightarrow\infty)$. This connection was pointed out by Nitanda and Suzuki, (2017); Mei et al., (2018); Chizat and Bach, (2018); Rotskoff and Vanden-Eijnden, (2022); Sirignano and Spiliopoulos, 2020b ; Sirignano and Spiliopoulos, 2020a in the case of $\lambda=0$, and used for showing the global convergence of the gradient flow for (1) by Mei et al., (2018); Chizat and Bach, (2018).
@@
$\mathrm{d}X_{t}=-\nabla\frac{\delta F(\mu_{t})}{\delta\mu}(X_{t})\mathrm{d}t+\sqrt{2\lambda}\mathrm{d}W_{t},~{}~{}~{}\mu_{t}=\mathrm{Law}(X_{t}),$ (2)
where $\{W_{t}\}_{t\geq 0}$ is the $d$-dimensional standard Brownian motion and $\nabla\frac{\delta F(\mu)}{\delta\mu}$ is the Wasserstein gradient that is the gradient of the first-variation $\frac{\delta F(\mu)}{\delta\mu}$ of $F$. Even though several optimization methods (Nitanda et al.,, 2021; Oko et al.,, 2022; Chen et al.,, 2023) that can efficiently solve the above problem with polynomial computational complexity have been proposed, MFLD remains an interesting research subject because of the above connection to the noisy gradient descent. In fact, recent studies showed that MFLD globally converges to the optimal solution (Hu et al.,, 2019; Mei et al.,, 2018) thanks to noise perturbation and that its convergence rate is exponential in continuous-time under uniform log-Sobolev inequality (Nitanda et al.,, 2022; Chizat,, 2022).
@@ 1.1 Contributions
Here, we mention the proof strategy of Theorem 1. Langevin dynamics (LD) is a special case of MFLD corresponding to the case where $F$ is a linear functional. It is well known that even with a single particle, we can simulate LD and the particle converges to the target Gibbs distribution. This means that the particle approximation error is due to the non-linearity of $F$. Therefore, in our analysis, we carefully treat the non-linearity of $F$ and obtain an expression for the particle approximation error using the Bregman divergence induced by $F$. Finally, we relate this divergence to the variance of an $N$-particle neural network and show the error of $O(1/N)$. This proof strategy is quite different from existing ones and is simple. Moreover, it leads to an improved approximation error as mentioned above. We refer the readers to Section 4 for details about the proof.
@@ 2.1 Problem setting
and say $G$ is convex when for $\forall\mu,\mu^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
@@
There exists $M_{1},M_{2}>0$ such that for any $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $x\in\mathbb{R}^{d}$, $\left|\nabla\frac{\delta F_{0}(\mu)}{\delta\mu}(x)\right|\leq M_{1}$ and for any $\mu,\mu^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $x,x^{\prime}\in\mathbb{R}^{d}$,
@@ 2.2 Mean-field Langevin dynamics and finite-particle approximation
where $\xi_{k}^{i}\sim\mathcal{N}(0,I_{d})~{}(i\in\{1,\ldots,N\})$ are i.i.d. standard normal random variables. By extending the proof techniques developed by Chen et al., (2022), Suzuki et al., 2023b proved the uniform-in-time propagation of chaos for MFLD (12); there exist constants $C_{1},C_{2}>0$ such that
@@
where $\mu^{(N)}_{k}=\mathrm{Law}(\mathbf{X}_{k})$. The last two terms are due to time-discretization and finite-particle approximation, respectively. The finite-particle approximation error $O(\frac{\lambda}{\alpha N})$ appearing in (11), (13) means the deterioration as $\alpha\rightarrow 0$. Considering typical estimation $\alpha\gtrsim\exp(-\Theta(1/\lambda))$ (e.g., Theorem 1 in Suzuki et al., 2023b ) of LSI-constant using Holley and Stroock argument (Holley and Stroock,, 1987) or Miclo’s trick (Bardet et al.,, 2018), these bounds imply that the required number of particles increases exponentially as $\lambda\rightarrow 0$.
@@ 3.1 LSI-constant free particle approximation error for mean-field neural networks
$\ell(\cdot,y)$ is convex and $L$-smooth, and $h(X,z)$ $(X\sim\mu_{*})$ has a finite-second moment;
@@
  • There exists $R>0$ such that for any $z\in\mathcal{Z}$, $\mathbb{E}_{X\sim\mu_{*}}[|h(X,z)|^{2}]\leq R^{2}$.
We can directly verify this assumption for mean-field neural networks using a bounded activation function (Nitanda et al.,, 2022; Chizat,, 2022; Chen et al.,, 2022; Suzuki et al., 2023b, ) and standard loss functions such as logistic loss and squared loss. The following is the main theorem that bounds $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})-\mathcal{L}(\mu_{*})$. The proof is deferred to Section 4 and Appendix 5.1.
@@
$\frac{\lambda}{N}\mathrm{KL}(\mu^{(N)}_{*}\|\mu^{\otimes N}_{*})\leq\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})-\mathcal{L}(\mu_{*})\leq\frac{LR^{2}}{2N}.$ (15)
@@ 3.2.1 Convergence of the mean-field Langevin dynamics
By setting $g=\sqrt{\frac{\mathrm{d}\mu^{(N)}}{\mathrm{d}\mu^{(N)}_{*}}}$, Assumption 4 leads to $\mathrm{KL}(\mu^{(N)}\|\mu^{(N)}_{*})\leq\frac{1}{2\bar{\alpha}}\mathbb{E}_{\mu^{(N)}}\left[\left\|\nabla\log\frac{\mathrm{d}\mu^{(N)}}{\mathrm{d}\mu^{(N)}_{*}}\right\|_{2}^{2}\right]$. For instance, using Holley and Stroock argument (Holley and Stroock,, 1987) under the boundedness assumption $|F_{0}(\mathbf{x})|\leq B~{}(\forall\mathbf{x}\in\mathbb{R}^{dN})$, we can verify LSI on $\mu^{(N)}_{*}$ with a constant $\bar{\alpha}$ that satisfies: $\bar{\alpha}\geq\frac{2\lambda^{\prime}}{\lambda}\exp\left(-\frac{4NB}{\lambda}\right)$. For the detail, see Appendix 6.
@@
    $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{t})-\mathcal{L}(\mu_{*})\leq\frac{LR^{2}}{2N}+\exp(-2\bar{\alpha}\lambda t)\left(\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{0})-\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})\right),$
@@
    $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{k})-\mathcal{L}(\mu_{*})\leq\frac{LR^{2}}{2N}+\frac{\delta^{(N)}_{\eta}}{2\bar{\alpha}\lambda}+\exp(-\bar{\alpha}\lambda\eta k)\left(\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{0})-\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})\right),$
@@
The term of $\frac{LR^{2}}{2N}$ is the particle approximation error inherited from Theorem 1. Again our result shows the LSI-constant independence particle approximation error for MFLD unlike existing results (Chen et al.,, 2022; Suzuki et al., 2023b, ) where their error bounds $O(\frac{\lambda}{\alpha N})$ scale inversely with LSI-constant $\alpha$ as seen in (11) and (13). Hence, the required number of particles to achieve $\epsilon$-accurate optimization: $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)})-\mathcal{L}(\mu_{*})\leq\epsilon$ suggested by our result and Chen et al., (2022); Suzuki et al., 2023b are $N=O(\frac{1}{\epsilon})$ and $N=O(\frac{\lambda}{\alpha\epsilon})$, respectively. Whereas the iterations complexity of MFLD (12) is $O(\frac{1}{\bar{\alpha}^{2}\lambda\epsilon}\log\frac{1}{\epsilon})$ which is same as that in Suzuki et al., 2023b up to a difference in LSI constants $\alpha$ or $\bar{\alpha}$.
@@ 3.2.2 Sampling guarantee for $\mu_{*}$
    $\lambda\mathrm{KL}(\mu^{(N)}_{t,1}\|\mu_{*})\leq\mathcal{L}(\mu^{(N)}_{t,1})-\mathcal{L}(\mu_{*})\leq\frac{LR^{2}}{2N}+\exp(-2\bar{\alpha}\lambda t)\Delta^{(N)}_{0},$
@@
    $\lambda\mathrm{KL}(\mu^{(N)}_{k,1}\|\mu_{*})\leq\mathcal{L}(\mu^{(N)}_{k,1})-\mathcal{L}(\mu_{*})\leq\frac{LR^{2}}{2N}+\frac{\delta^{(N)}_{\eta}}{2\bar{\alpha}\lambda}+\exp(-\bar{\alpha}\lambda\eta k)\Delta^{(N)}_{0}.$
@@
For any distribution $\mu^{(N)}\in\mathcal{P}_{2}(\mathbb{R}^{dN})$ whose marginal $\mu^{(N)}_{i}$ on $i$-th coordinate $x^{i}$ ($i\in\{1,\ldots,N\}$) are identical to each other, it follows that by the convexity of the objective function and the entropy sandwich (Nitanda et al.,, 2022; Chizat,, 2022): $\lambda\mathrm{KL}(\mu\|\mu_{*})\leq\mathcal{L}(\mu)-\mathcal{L}(\mu_{*})$ $(\forall\mu\in\mathcal{P}_{2}(\mathbb{R}^{d}))$,
@@ 3.2.3 Uniform-in-time Wasserstein propagation of chaos
    $\frac{1}{N}W_{2}^{2}(\mu^{(N)}_{t},\mu^{\otimes N}_{t})\leq\frac{4}{\alpha\lambda}\left(\frac{LR^{2}}{2N}+\exp(-2\bar{\alpha}\lambda t)\Delta^{(N)}_{0}+\exp(-2\alpha\lambda t)\Delta_{0}\right).$
@@
    $\frac{1}{N}W_{2}^{2}(\mu^{(N)}_{k},\mu^{\otimes N}_{k})\leq\frac{4}{\alpha\lambda}\left(\frac{LR^{2}}{2N}+\frac{\delta^{(N)}_{\eta}}{2\bar{\alpha}\lambda}+\frac{\delta_{\eta}}{2\alpha\lambda}+\exp(-\bar{\alpha}\lambda\eta k)\Delta^{(N)}_{0}+\exp(-2\alpha\lambda\eta k)\Delta_{0}\right),$
@@
Applying the convergence rates of finite- and infinite-particle MFLDs (Theorem 2 and Nitanda et al., (2022)), we conclude the proof. For completeness, we include the auxiliary results used in the proof in Appendix 6. ∎
Corollary 2 uniformly controls the gap between $N$-particle system $\mu^{(N)}_{t}/\mu^{(N)}_{k}$ and its mean-field limit $\mu^{\otimes N}_{t}/\mu^{\otimes N}_{k}$. Again this result shows an improved particle approximation error $O(\frac{1}{\alpha\lambda N})$ over $O(\frac{1}{\alpha^{2}N})$ (Chen et al.,, 2022; Suzuki et al., 2023b, ). Additionally, the propagation of chaos result in terms of TV-norm can be proven by using Pinsker’s inequality instead of Talagrand’s inequality in the proof. For the continuous-time MFLDs, we get
$\frac{1}{N}\mathrm{TV}^{2}(\mu^{(N)}_{t},\mu^{\otimes N}_{t})\leq\frac{1}{\lambda}\left(\frac{LR^{2}}{2N}+\exp(-2\bar{\alpha}\lambda t)\Delta^{(N)}_{0}+\exp(-2\alpha\lambda t)\Delta_{0}\right),$
and TV-norm counterpart for the discrete-time can be derived similary.
@@
In this section, we provide the proof sketch of Theorem 1. Our analysis carefully treats the non-linearity of $F_{0}$ because the particle approximation errors, the gap between $\mathcal{L}/\mu_{*}$ and $\mathcal{L}^{(N)}/\mu^{(N)}_{*}$, come from this non-linearity. In fact if $F_{0}$ is a linear functional: $F_{0}(\mu)=\mathbb{E}_{\mu}[f]$ ($\exists f:\mathbb{R}^{d}\rightarrow\mathbb{R}$), then $\mathcal{L}^{(N)}(\mu^{(N)})=\sum_{i=1}^{N}\mathbb{E}_{X^{i}\sim\mu^{(N)}_{i}}[f(X^{i})+\lambda^{\prime}\|X^{i}\|_{2}^{2}]+\lambda\mathrm{Ent}(\mu^{(N)})\geq\sum_{i=1}^{N}\mathcal{L}(\mu_{i}^{(N)})$, where $\mu^{(N)}_{i}$ are marginal distributions on $X^{i}$. This results in $\mu^{(N)}_{*}=\mu^{\otimes N}_{*}$ and $\mathcal{L}^{(N)}(\mu^{(N)}_{*})=N\mathcal{L}(\mu_{*})$, and thus there is no approximation error by using finite-particles as also deduced from Theorem 1 with $L=0$. Therefore, we should take into account the non-linearity of $F_{0}$ to tightly evaluate the gap between $\mathcal{L}$ and $\mathcal{L}^{(N)}$.
@@
$=\exp\left(-\frac{N}{\lambda}\left(F(\mu)+\left\langle\frac{\delta F(\mu)}{\delta\mu},\mu_{\mathbf{x}}-\mu\right\rangle+B_{F}(\mu_{\mathbf{x}},\mu)\right)\right)$
$\propto\exp\left(-\frac{N}{\lambda}B_{F}(\mu_{\mathbf{x}},\mu)\right)\prod_{i=1}^{N}\exp\left(-\frac{1}{\lambda}\frac{\delta F(\mu)}{\delta\mu}(x^{i})\right)$
@@
Intuitively, $\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[B_{F}(\mu_{\mathbf{X}},\mu_{*})\right]$ is small because the empirical distribution $\mu_{\mathbf{X}}$ ($\mathbf{X}\sim\mu^{\otimes N}_{*}$) converges to $\mu_{*}$ by law of large numbers. Indeed, more simply, we can relate this term to the variance of an $N$-particle mean-field model $h_{\mu_{\mathbf{X}}}(z)$ ($\mathbf{X}\sim\mu_{*}^{\otimes N}$), yielding a bound: $\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[B_{F}(\mu_{\mathbf{X}},\mu_{*})\right]\leq\frac{LR^{2}}{2N}$ (see the proof of Theorem 1 in Appendix 5). Then, we arrive at Theorem 1.